The following information applies to the next two questions. Dée Trader opens a brokerage account and purchases 300 shares of Internet Dreams at $40 per share. She borrows $4,000 from her broker to help pay for the purchase. Question 1 (1 point) What is the initial margin in Dées account when she first purchases the stock? Enter your answer as a decimal, rounded to two decimal places Your Answer Answer uestion 2 (1 point) If the share price falls to $30 per share, what is the margin in her account? Enter your answer as a decimal, rounded to two decimal places. Your Answer: Answer
Answer #1

Answer for question 1

Stock is purchased for = 300 * 40 = $ 12000

Amount borrowed from broker = $4000

Therefore initial margin when she purchased the stock = 12000-4000 = $8000

Answer for question 2

If stock price falls to $30 value of stock = $30 * 300 = 9000

Therefore margin in the account if price falls to $30 = 9000-4000 = $5000